OpenStudy (anonymous):

3. A local store owner pays her employees time-and-a-half for overtime. That means if an employee works more than 40 hours per week, the store will pay him 1.5 times his regular hourly wage of $6.50. (10 points) (a) Write and graph a piecewise function that illustrates an employee's weekly pay, W, as a function of the number of hours, h, that employee works.

OpenStudy (anonymous):

w= 40 * 6.50 + (h-40) * 1.5 * 6.50 ............ 40 * 6.50 for the required 40 hrs and (h-40) for any extra hours w= 9.75h -390 + 260 w= 9.75 h -130 i think this is the function.